A car enters the highway at point a and travels west at a constant speed of 55 miles per hour. One hour later, another car enters the highway at point a and travels west at a constant speed of 65 miles per hour. How long does it take the second car to overtake the first?
Answer by richwmiller(9135)

55*(t+1)=65t
55t+55=65t
10t=55
t=55/10=5.5 hrs